At a local college, of the 162 male students are smokers and 648 are non-smokers. Of the female students, 140 are smokers and 560 are non-smokers. A male student and a female student from the college are randomly selected for a survey. What is the probability that both are smokers? Do not round your answer. (If necessary, consult a list of formulas.)
Answer #1

Answer)

Probability is given by favorable/total

Probability that male student is smoker = 162/(162+648) = 0.2

Probability that female student is smoker = 140/(140+560) = 0.2

Probability that both are smoker = 0.2*0.2 = 0.04